How do i solve this
(x+2) squared =3(x+2)

Answer by Theo(13342)

(x+2)^2 = x^2 + 4x + 4

3*(x+2) = 3x + 6

your equation of (x+2)^2 = 3(x+2) becomes x^2 + 4x + 4 = 3x + 6

subtract 3x and 6 from both sides of the equation to get x^2 + 4x + 4 - 3x - 6 = 0

combine like terms to get x^2 + x - 2 = 0

factor this equation to get (x-1) * (x+2) = 0

solve for x to get -2 or x = 1.

those are your solutions.

to confirm they are applicable to the problem, replace x with 2 and then replace x with -1 in the original equations.

the original equation is (x+2)^2 = 3(x+2)

when x = -2, this equation becomes 0 = 0.

when x = 1, this equation becomes 9 = 9.

this confirms both solutions are correct.